I am having problem this this question could you please help !Five apples and a pear cost $ 3.93. If an apple costs $ 0.33, how much does a pear cost?
Answer by jim_thompson5910(35256)

a = cost of one apple in dollars
p = cost of one pear in dollars

an apple costs $ 0.33 is given information, so a = 0.33. The value of p is unknown

------------------------------------------------------------

"Five apples and a pear cost $ 3.93" so we can say

5a + 1p = 3.93

plug in a = 0.33 and solve for p

------------------------------------------------------------

5a + 1p = 3.93

5*0.33 + 1p = 3.93 Replace every copy of 'a' with '0.33'

1.65 + 1p = 3.93

1.65 + 1p - 1.65 = 3.93 - 1.65 Subtract 1.65 from both sides

1p + 0 = 2.28

1p = 2.28

p = 2.28

Therefore, one pear costs $2.28